OpenStudy (anonymous):

Simplify (1-x)^-2*(x-1)

OpenStudy (anonymous):

I don't know what to do with a negative exponent

OpenStudy (anonymous):

You change the position in a rational expression. Rational expressions are just fractions. A negative exponent just "swaps" and becomes positive. So, 1/2^-2 , the 2^-2 power is in the denominator. It still keeps the exponent of ^2, but gets swapped to the numerator.

OpenStudy (anonymous):

Example: 1 / 2^-2 is 2^2/1 or 4/1 or 4. The exponent stays the same, you just switch the whole thing from the numerator to denominator, or vise versa.

OpenStudy (anonymous):

\[\frac{ 1 }{ 2^{-2} }\] swaps places to \[\frac{ 2^{2} }{ 1 }\] but make sure you understand that the "1" is implied. Anything by itself is the same thing as being over one. The one didn't switch places. Only what had the negative exponent did.

OpenStudy (anonymous):

would have been the same thing as \[1\times 2^{2} \] but we don't keep the one, because anything multiplied by one is itself, so we can just get rid of it. Same thing as anything by itself is the same thign as dividing by one, so we get rid of that too.